OpenStudy (anonymous):

WHAT ARE THE 3 PLOTTING POINT FOR -3Y=-X-7

OpenStudy (anonymous):

rewrite as 3y=x+7 or y= x/3 +7/3 put x=0,1,2 to get three points on line

OpenStudy (anonymous):

K I'LL TRY IT

OpenStudy (anonymous):

should work i guess

OpenStudy (anonymous):

IS FIRST ONE 7,3

OpenStudy (anonymous):

u mean (0, 7/3) ???

OpenStudy (anonymous):

IM CONFUSED

OpenStudy (anonymous):

substitute x=0 in y= x/3 +7/3 , u get y=0 +7/3=7/3 hence u have got one point (0,7/3)

OpenStudy (anonymous):

what r other people's views on this?

OpenStudy (anonymous):

2 ND POINT I GOT 8,3 BUT ON THE 3RD ONE IS IT 3,0 OR 3,1 LOST ON LAST ONE

OpenStudy (anonymous):

second point could be(8,5)

OpenStudy (anonymous):

OK